# Look the Other Way: Designing ‘Positive’ Molecules with Negative Data via Task Arithmetic

### v1

@@ 4.2 Zero-Shot Molecule Design
We next explore the capabilities of molecular task arithmetic for dual-objective de novo design. This objective is central in drug discovery, for instance, for multi-target drug design, or selectivity optimization. We define three dual-objective tasks via the molecular properties above: (a) high fraction of sp3-hybridized carbons and high number of hydrogen bond donors, (b) high lipophilicity and low TPSA, and (c) high number of rings and low TPSA. These pairs of properties were selected to be minimally correlated (absolute correlation lower than 0.3; Table A3), so that the models neither face a conflicting nor a trivial task.
To apply molecular task arithmetic in a dual-objective setting (Fig. 1d), we sum the scaled vectors yielding the best performance for each single task (Eq. 3) and scale the final vector in increasing $\lambda$ values (0.05 to 1.00 with a step size of 0.05). For finetuning, we finetuned the models on the training set of the single tasks sequentially (using positive molecules), as in recent literature 56, 57, with both task permutations. We reported the number of (successful) design clusters, success rate, and validity for the $\lambda$ and permutation yielding the highest number of successful design clusters for molecular task arithmetic and finetuning, respectively.
Fig. 4: Zero-shot dual objective molecule design. Models are trained on randomized SMILES representation with sequential supervised finetuning and molecular task arithmetic to design molecules that possess two task properties simultaneously. (a) Number of successful design clusters, (b) number of clusters, (c) success rate, and (d) validity are measured. Experiments are repeated five times, and the mean (bar height) and standard deviation (error bar) are reported.
@@
4.3 Effect of Task Vector Scaling
@@
Fig. 5: Impact of molecular task vector scaling. For each design task defined on randomized SMILES strings, molecular task arithmetic is applied in 19 increasing scaling factors ($\lambda$; Eq. 3). 100,000 molecules are designed with all values of $\lambda$ and validity and success rate are measured across five training splits. Lines denote the means across runs, and shaded areas describe the standard deviations.
@@ 4.4 Boosting Few-Shot Molecule Design
Fig. 6: Few-shot molecule design with molecular task arithmetic. Models are trained on an increasing number of positive data points with supervised finetuning and molecular task arithmetic. Molecular task arithmetic is first applied using negative data, and then the resulting model is fine-tuned with the available task data. The mean and the standard deviation of the number of successful design clusters across five splits are computed (lines and shaded regions, respectively).
@@
We next delve into another interesting problem in drug discovery: few-shot molecule design. In few-shot molecule design, some positive labeled molecules are already available, which are used for molecule design. Additionally, negative molecules are usually available (and in general more abundant), but are not leveraged in traditional finetuning pipelines. Here we set out to investigate whether molecular task arithmetic can allow leveraging negative molecules alongside positive ones to boost the capacity of models to generate molecules with desirable properties.
@@
We applied molecular task arithmetic as in the zero-shot design experiments (Eq. 3) first, and then fine-tunex the obtained model with the known positive molecules (Fig. 1e). In other words, this pipeline replaces the pretraining model used for supervised finetuning with the molecular task arithmetic model, to harness the benefits of abundant and structurally diverse negative molecules. We compare the proposed approach to traditional supervised finetuning by computing the number of design clusters when increasing the number of positive molecules.
Across all 20 tasks and the number of positive molecules studied, molecular task arithmetic surpassed the number of successful clusters obtained with finetuning, up to a 10K increase (Fig. 6). A deeper look reveals that molecular task arithmetic can design at least as many successful designs as supervised finetuning (Fig. A8), and yet maintain the design diversity (Fig. A9). Overall, these experiments point to a synergy: whenever positive molecules are available, molecular task arithmetic can be combined with supervised finetuning to increase the number of diverse hits.
@@
Molecular task arithmetic can reach maximum performance even with fewer than 50 positive molecules. The number of clusters gradually decreases as the model is finetuned with further molecules. Such performance fluctuations, however, have a smaller magnitude than vanilla finetuning, underlining the robustness of molecular task arithmetic. Taken together, the experiments highlight molecular task arithmetic as a data-efficient, robust, and well-performing transfer learning strategy for few-shot molecule design. Combined with its simplicity, molecular task arithmetic can potentially become the new standard for transfer learning in de novo design, by allowing one to leverage both positive and negative data simultaneously.
5 Related Work
Task arithmetic.
Discovering semantic directions in the model weight space has emerged as a new research direction that offers opportunities to edit model behavior with simple arithmetic operations 25, 27, 28, 29. Named as task arithmetic, this perspective has enabled higher performance with model merging 26, in-context task learning28, and zero-shot information retrieval27. Its capability to recognize task directions from finetuning of opposite labels offers potential for molecular field, where negative samples are dominantly more available than positive ones. Here we explored this potential.
Molecule design.
Deep learning for de novo drug design has been an active field in the last decade 42, 58, 59. Early approaches trained sequence models on string representations of molecules with a next-token prediction task, using language modeling42, 58, 59. Recent works used graph and 3D-based representations to represent molecules and graph neural networks and diffusion for learning 60, 61, 62, 63, 64, 65. While these approaches enable more fine-grained representation and high steerability in learning, decoder-only language models of molecular strings, termed chemical language models, remain popular to date due to their simplicity and performance40. Furthermore, one molecule can have multiple SMILES strings, which enables data augmentation for deep learning in low data 34, 56, 66, 45.
6 Limitations
Molecular task arithmetic relies on negative data. While negative molecules are abundant due to the characteristics of the chemical space, this still forms a limitation to applying molecular task arithmetic. Our experiments also revealed that the validity of generations drops with molecular task arithmetic. Smarter weight edits29 and/or alternative architectures should be studied to minimize this side effect. Last, this study tested molecular task arithmetic only for de novo design using LSTMs on SMILES strings. Experiments on other molecular generation tasks with more architectures and representations are needed to determine the limits of molecular task arithmetic as a new transfer learning strategy for deep learning in chemistry.
In this study, molecular task arithmetic was tested on molecular properties that can be exactly calculated and that vary linearly. It is yet to be determined whether negating more complex and non necessarily linear properties (e.g., inactivity, in the case of ligand-protein interaction) would yield molecules with the desired properties (e.g., biological activity).
@@ 7 Conclusion
Drug discovery datasets are usually small and lack structural diversity. While deep generative modeling and transfer learning have already been transformative for drug design, they still must face this inherent challenge. This work proposed a remedy: molecular task arithmetic. Molecular task arithmetic is a new transfer learning strategy that leverages the diverse and abundant negative molecules. Our approach unlocks zero-shot ligand-based molecule design in single- and dual-objective settings – creating a new paradigm for drug discovery. Experiments across ten design tasks show that molecular task arithmetic – which uses no positive examples – can design more diverse hits that fulfill desirable properties than models trained solely on positive data. We show that molecular task arithmetic can also leverage the available positive molecules to further propel chemical space exploration – outperforming traditional transfer learning across all design tasks. Thanks to its simplicity and performance, molecular task arithmetic bears the potential to become a well-established transfer learning strategy for drug discovery, a field where negative data abound.
@@
A chemical language model with an LSTM backbone is pretrained on a previously curated dataset of 1.5M drug-like molecules from ChEMBL[21, 67]. 100 models were trained for either canonical or randomized SMILES strings, with random hyperparameters (Table A1). For each representation, the model that generates the highest number of unique and new designs was selected for the finetuning stage.
The finetuning sets were curated from a previous study [24] by ensuring that no finetuning molecule is present in the pretraining set. The thresholds for high and low values are computed by rounding the median of the molecular property values among the pretraining molecules (Table A2 and Fig. A1). Five molecule sets are curated for each task (1024 train, 256 validation, 256 test molecules per split), where each molecule carries the task property. For models trained on randomized SMILES, 10-fold SMILES augmentation is used during finetuning. Early stopping on validation loss (cross-entropy) with a patience of five epochs and loss tolerance of $1\times 10^{-5}$ is employed. 100,000 molecules were generated with each model.
@@
  • Number of clusters: the number of clusters identified by the sphere exclusion algorithm. Tanimoto distance on extended connectivity fingerprints [68] is used with a threshold of 65% to identify the number of distinct clusters. The metric is linked to #Circles [69], and the rdkit implementation is used. The higher the number of clusters, the more diverse the design library – a desired trait for drug discovery [70].
@@
  • Number of successful molecule clusters: the number of cluster centers that possess the task property. This metric combines the evaluation of diversity and accuracy, and is used as the main evaluation metric across the study. Higher values are preferred.
@@
Fig. A4: Side effects of conditioning strategies and pretraining. Compuations are described in (Fig. 3). Bar heights display the means across five training splits, and error bars display the standard deviations. In 32 of 40 comparisons, molecular task arithmetic caused less side-effects.
@@
Property No. H. Donor No. Rings logP TPSA Frac. sp3 C
No. H. Donor - -0.09 -0.28 0.17 0.02
No. Rings -0.09 - 0.40 0.26 0.13
logP -0.28 0.40 - -0.12 0.08
TPSA 0.17 0.26 -0.12 - 0.57
Frac. sp3 C 0.02 0.13 0.08 0.57 -
@@
Fig. A7: Impact of molecular task vector scaling on canonical SMILES. For each design task defined on canonical SMILES strings, molecular task arithmetic is applied in increasing scaling factors ($\lambda$; Eq. 2). 100,000 molecules are designed with all $\lambda$s and validity and success rate are measured across five training splits. Lines denote the means across runs, and shaded areas describe the standard deviations.

### v2

@@ 5.2 Zero-Shot Molecule Design
We next explore the capabilities of molecular task arithmetic for zero-shot dual-objective de novo design. This objective is central in drug discovery, for instance, for multi-target drug design, or selectivity optimization. To apply MTA in a dual-objective setting (Fig. 1b), we sum the scaled vectors yielding the best performance for each single task (Eq. 3) and scale the final vector in increasing $\lambda$ values (0.05 to 1.00 with a step size of 0.05). For finetuning, we finetuned the models on the training set of the single tasks sequentially (using positive molecules), as in recent literature ballarotto2023lowdata, with both task permutations. We reported the scores for the $\lambda$ and permutation yielding the highest number of successful design clusters for MTA and finetuning.
Fig. 4: Zero-shot dual objective molecule design. Models are trained on randomized SMILES representation with sequential supervised finetuning and molecular task arithmetic to design molecules that possess two task properties simultaneously (average and standard deviation across five splits).
@@
5.3 Effect of Task Vector Scaling and Practical Implications
@@
Fig. 5: Task vector scaling, across 19 increasing scaling factors ($\lambda$; Eq. 3). For each $\lambda$, validity and success rate for 100,000 designs were computed (average and standard deviation; five training splits).
@@ 5.4 Boosting Few-Shot Molecule Design
Fig. 6: Few-shot molecule design. Models are trained on an increasing number of positive molecules. Molecular task arithmetic (MTA) was first applied using negative data, and then the resulting model was finetuned with the available task data. Average and standard deviation across five splits reported.
@@
We next delve into another interesting problem in drug discovery: few-shot molecule design. In few-shot molecule design, some positive labeled molecules are already available, which are used for molecule design. Additionally, negative molecules are usually available (and in general more abundant), but are not leveraged in traditional finetuning pipelines. Here we set out to investigate whether molecular task arithmetic (MTA) can allow leveraging negative molecules alongside positive ones to boost the capacity of models to generate molecules with desirable properties.
@@
We applied MTA as in the zero-shot design experiments (Eq. 3) first, and then fine-tuned the obtained model with the known positive molecules (Fig. 1c). In other words, this pipeline replaces the pretraining model used for supervised finetuning with the MTA model, to harness the benefits of abundant and structurally diverse negative molecules. We compare the proposed approach to traditional supervised finetuning by computing the number of design clusters when increasing the number of positive molecules.
@@
Across 20 tasks and the number of positive molecules studied, MTA surpassed the number of successful clusters obtained with finetuning, up to a 10K increase (Fig. 6). A deeper look reveals that MTA can design at least as many successful designs as supervised finetuning (Fig. A12), and yet maintain the design diversity (Fig. A13). Overall, these experiments point to a synergy: whenever positive molecules are available, MTA can be combined with supervised finetuning to increase the number of diverse hits. MTA can reach maximum performance even with fewer than 50 positive molecules. The number of clusters gradually decreases as the model is finetuned with further molecules. Such performance fluctuations, however, have a smaller magnitude than vanilla finetuning, underlining the robustness of MTA. Taken together, the experiments highlight MTA as a data-efficient, robust, and well-performing transfer learning strategy for few-shot molecule design. Combined with its simplicity, MTA can potentially become the new standard for transfer learning in de novo design, by allowing one to leverage both positive and negative data simultaneously.
5.5 Bioactive Molecule Design with Inactivity Data
Fig. 7: Bioactive molecule design. Models are trained with well- and poor-docking molecules. 10,000 designs are generated with each strategy and 10,000 pretraining molecules are included as a control. Number of successful clusters (a, c) and success rate (b, d) are computed for each set.
We now focus on designing molecules possessing a more complex property: a good docking score on a chosen protein target, starting from molecules that are not interacting with the target. Since ligand–protein interaction depends on shape complementarity, chemical compatibility, and binding-site dynamics, this setting presents a particularly challenging test case. We used Autodock-GPU eberhardt2021autodock for docking, and docked 100,000 randomly selected pretraining molecules against F2, NR3C1, and PPARD. We found that well-docking molecules constitute less than 5% for each target (Fig. A3), highlighting the scarcity of positive molecules and the difficulty of this task.
We evaluated molecular task arithmetic (MTA) to design well-docking molecules in zero- and few-shot settings. In the zero-shot setting, MTA yielded 2.3-2.8 times more successful molecule clusters compared to the pretraining set (10,000 control molecules) (Fig. 7a), with better docking scores (2.1-2.5 fold improvement, Fig. 7b). In the few-shot setting, MTA created 643 to 1100 more successful design clusters than traditional finetuning (Fig. 7c), and its designs were structurally more diverse from the training set (Fig. A14), suggesting that task arithmetic can create a more diverse set of promising molecular candidates. Overall, fine-tuning shows a higher success rate (7.29% and 3.84% higher for F2 and NR3C1, respectively), whereas for PPARD, MTA boosts the success rate by 5.18%. Notably, PPARD has the smallest number of positive molecules (12 molecules; 13% of F2 and 9% of NR3C1; Table A3). Taken together, these results demonstrate that MTA can design molecules with complex, desirable properties without any such labeled data points (Fig. A15). Furthermore, MTA can increase the diversity of successful designs in general, with particular promise in low data settings. These findings further demonstrate the potential of task arithmetic to mitigate data availability bottlenecks and open exciting new frontiers for de novo molecule design.
5.6 Molecular Task Arithmetic for Protein Design with Transformers
To extend our findings beyond LSTMs and small molecules, we applied molecular task arithmetic (MTA) to a large pretrained transformer language model for protein design. We experimented with ProtGPT2 ferruz2022protgpt2, a 738M-parameter model pretrained on amino acid sequences, to design proteins with more structurally-ordered regions. A high degree of intrinsic disorder in proteins leads to multiple possible 3D conformations, posing a substantial challenge for designing proteins with a specific fold listov2024opportunities. We curated proteins lacking well-defined 3D structures from DisProt aspromonte2024disprot as highly disordered (negative) examples and applied MTA. We benchmarked our approach against ProtGPT2 finetuned on a highly structured subset of the pretraining set.
Table 1: Ordered amino acid ratio, globular protein ratio, and redundancy across datasets and design methods. Best (bold) and second-best (underline) approaches are highlighted.
Method Ordered AA Ratio Globular Ratio Redundancy (@30%)
Pretraining Set 83% 88% 10%
Finetuning Set 97% 100% 7%
DisProt 28% 0% 44%
ProtGPT2 80% 88% 12%
Finetuning 97% 100% 35%
MTA 90% 92% 22%
The fraction of amino acids predicted to be ordered by IUPred3 erdHos2021iupred3 increases from 80% to 90% with MTA (Table 1), and the fraction of designs with at least one globular domain increases from 88% to 92%, compared to the pretrained model. Finetuning on highly structured proteins yields 97% and 100% on these metrics, respectively, at the cost of increased redundancy. Finetuning designs have a redundancy of 35% at 30% identity, meaning that 35% of the designed sequences are similar to at least one other finetuning design over 30%. On the contrary, MTA exhibits lower redundancy values (22%), suggesting that this approach can design more diverse proteins. DisProt (training set of MTA) is the most redundant library in the comparison; yet, MTA enables exploring high-diversity regions. Our experiments corroborate the versatility and capacity of MTA to achieve accuracy and diversity also with this additional modality, and model architecture and scale.
5.7 On Failure Modes of Molecular Task Arithmetic
A key parameter in MTA is the magnitude of the task subtraction, controlled by $\lambda$ (Eq. 3). Excessive values lead to a deterioration of SMILES validity (Fig. 5) and a corresponding drop in success rates. This indicates that “jumping too far” from the pretrained model can induce model collapse. We observe a strong correlation between validity and success rate, suggesting that validity can serve as a practical proxy for $\lambda$ selection when success rate is difficult to measure. ‘Selective task arithmetic’ bowen2024beyond (i.e., updating certain weights only), or alternative model architectures and representations could be explored to mitigate the validity drop. Moreover, MTA could extrapolate beyond the training data, albeit depending on the design task and the frequency of positive molecules in the pretraining set. This means that the extrapolation capacities will have to be analyzed on a case-by-case basis – posing potential constraints for properties that are difficult to compute. Finally, while MTA showed promising results for dual-objective design, combining task vectors of conflicting tasks might introduce task collapse, constraining its applicability in multi-objective settings.
@@ 6 Conclusion
We proposed a transfer learning strategy to sidestep the low-data bottleneck in drug discovery: molecular task arithmetic (MTA). MTA leverages the diversity and abundance of negative molecules and can design molecules zero-shot. Experiments across architectures, molecular entities, and task difficulties show that MTA can design more diverse hits than models trained on positive data. While we studied molecular task arithmetic as an alternative (and complementary) approach to supervised finetuning, preliminary results highlight its integration with goal-directed optimization using reinforcement learning as an exciting research direction (see Appendix and Table A6). Thanks to its simplicity and performance, we expect MTA to attract further research and become a prominent transfer learning strategy for drug discovery, a field where negative data is abundant.
@@ Acknowledgements
SMILES strings
SMILES [weininger1988smiles] is the most popular string representation for chemical language modeling (Fig. A1a). They annotate the bonds, atoms, and branches in a molecular graph, starting from any non-hydrogen atom. This characteristic allows creating multiple SMILES strings per molecule, randomized SMILES, enabling data augmentation [bjerrum2017smiles]. Canonicalization algorithms were proposed to obtain a single SMILES string from any molecule, aiding in duplicate detection [brown2019guacamol]. Both SMILES formats are commonly used for molecule design [grisoni2023chemical], with randomized SMILES strings yielding higher chemical space exploration [arus2019randomized].
For modeling purposes, the SMILES string represents a molecule $m$ as $s_{1},\dots,s_{n}$, where $s_{i}$ is the $i^{th}$ token in the molecular string. The next token prediction language modeling objective is then defined as: $\max\sum_{m}^{M}\sum_{i}\log P(s_{i}|s_{<i};\theta)$, where $\theta$ is typically a long-short term memory network [segler2018generating] or a transformer [bagal2021molgpt], and $M$ is a dataset of SMILES strings (Fig. 1b).
Fig. A1: Molecular task arithmetic. (a) SMILES strings annotate atoms by periodic table symbols and use additional tokens for bonds, branches, and rings. (b) Chemical language modeling casts SMILES generation as next-token prediction (e.g., using a sequence model, such as an LSTM).
@@
A chemical language model with an LSTM backbone is pretrained on a previously curated dataset of 1.5M drug-like molecules from [gaulton2017chembl, ozccelik2024jungle]. 100 models were trained for either canonical or randomized SMILES strings, with random hyperparameters (Table A4). For each representation, the model that generates the highest number of unique and new designs was selected for the finetuning stage.
The finetuning sets were curated from a previous study [sun2017excape] by ensuring that no finetuning molecule is present in the pretraining set. The thresholds for high and low values are computed by rounding the median of the molecular property values among the pretraining molecules (Table A5 and Fig. A2). Five molecule sets are curated for each task (1024 train, 256 validation, 256 test molecules per split), where each molecule carries the task property. For models trained on randomized SMILES, 10-fold SMILES augmentation is used during finetuning. Early stopping on validation loss (cross-entropy) with a patience of five epochs and loss tolerance of $1\times 10^{-5}$ is employed. 100,000 molecules were generated with each model.
@@
  • Number of clusters: the number of clusters identified by the sphere exclusion algorithm. Tanimoto distance on extended connectivity fingerprints [rogers2010extended] is used with a threshold of 65% to identify the number of distinct clusters as suggested in previous work [xiemuch]. The metric is linked to #Circles [xiemuch], and the rdkit implementation is used. The higher the number of clusters, the more diverse the design library – a desired trait for drug discovery [renz2024diverse].
@@
  • Number of successful molecule clusters: the number of cluster centers that possess the task property. Successful designs are first extracted from the full list and then clustered to identify the structurally distant molecules. This metric combines the evaluation of diversity and accuracy, and is used as the main evaluation metric across the study. Higher values are preferred.
Docking.
We selected coagulation factor II (F2), glucocorticoid receptor (NR3C1), and Peroxisome proliferator-activated receptor delta (PPARD) proteins for docking experiments since they are studied in drug discovery contexts and docking yields high enrichment for these targets [garcia2022dockstring]. The binding site annotations, binding ligands, and non-binding ligands are curated from previous work [garcia2022dockstring]. Any ligand that contains atoms besides C, H, O, N, S, P, F, Cl, Br, I, or is present in the pretraining set is filtered out. Ligands with 10-40 non-hydrogen atoms and a canonical SMILES length below 80 are extracted. Salts, charges, and stereochemistry annotations are removed. Meeko is used to preprocess the proteins and ligands for docking. Autogrid4 is used to create interaction maps of the protein binding sites, and Autodock-GPU is used to run the docking simulations [santos2021accelerating].
20 docking simulations are run for each ligand, and the lowest binding energy across the runs is used as the docking score. 100,000 pretraining molecules are randomly sampled and also docked against each target using the same pipeline. The rounded values that define the lowest and highest 5% of the pretraining docking score distributions are used as thresholds to define ‘well-docking’ and ‘poor-docking’ molecules (Fig. A3). The thresholds for F2 are found to be -10.75 and -5.75 for well-docked and poorly docked molecules, respectively. For NR3C1 and PPARD; the same thresholds are computed as -12.00 and -6.00.
The actives with a docking score below the well-docking threshold, and inactives with a docking score above the poor-docking threshold are extracted for supervised finetuning and molecular task arithmetic. The molecules in each set are represented with randomized SMILES and augmented 10 times. The models are trained with early stopping with configurations the same as above. 10,000 molecules are generated with each trained model, and novel and unique designs are docked against the corresponding target using the same docking pipeline used for datasets.
We train models with molecular task arithmetic ($0.20<\lambda<0.90$) using poor-docking molecules for zero-shot design. For the few-shot setting, we apply supervised finetuning using positive molecules and task arithmetic as previously (Section 5.4). We produce 10,000 samples with each strategy and dock the novel and unique designs. We randomly sample 10,000 molecules from the pretraining set as a control and compute the number of successful clusters and success rate.
Side effects
A desirable trait of model conditioning is the possibility of having minimal side effects on the ‘off-target’ properties. Being able to minimize side effects would allow for a more precise steering of the model in the chemical space, by preserving desirable molecular properties (i.e., synthesizability, lack of toxicity, etc.). We quantify the side effects of both finetuning and molecular task arithmetic for all design tasks, on the remaining four descriptors. These descriptors have a low correlation with the ‘target’ properties (Table A1). For each task, we computed the Kolmogorov-Smirnov (KS) distance between the off-target molecular properties of the designs and the pretraining set (the lower the KS, the more similar the distributions). The usefulness of this analysis is corroborated by observing the KS distances of the designs obtained by pretraining (Fig. A7), which are all below 5%.
Across the ten design tasks for randomized SMILES, molecular task arithmetic showed a smaller mean distributional distance to the pretraining set in 32 of the 40 comparisons (Fig. A6), of which 15 were confirmed by a statistical test (Mann-Whitney U test, p-value << 0.01). On canonical SMILES strings, all models have a higher average distance (Fig. A8), possibly because canonicalization restricts the models to explore ‘side-effect-free’ portions of the chemical space. Overall, the analysis shows that molecular task arithmetic can combine its accuracy and diversity on the design task, with less side-effect than the traditional finetuning approach. This suggests that task arithmetic allows for a more controlled steering of the models in the desirable portions of the chemical space.
Molecular Task Arithmetic and Reinforcement Learning
We have run preliminary experiments to test the potential of molecular task arithmetic within goal-directed optimization with reinforcement learning. We have run five REINVENT loops with default parameters to design molecules with high and low TPSA values using pretrained and MTA models from the out-of-distribution design experiments (Section 5.2). With 100,000 molecules designed per run (Table A6), REINVENT with MTA obtained 15K and 35K more successful design clusters for low-tpsa and high-tpsa design tasks, respectively, than running REINVENT with the pretrained model. MTA also increased the success rate by 27% and 47%, while pretrained model yield 4K and 1K more design clusters in the respective tasks. While a systematic, large-scale analysis across optimization algorithms and design tasks is needed to accurately assess the harmony between MTA and RL, this preliminary analysis shows that MTA can also be used to kick-start an RL loop for zero-shot design with external reward functions.
Property No. H. Donors No. Rings logP TPSA Frac. sp3 C
No. H. Donors - -0.09 -0.28 0.17 0.02
No. Rings -0.09 - 0.40 0.26 0.13
logP -0.28 0.40 - -0.12 0.08
TPSA 0.17 0.26 -0.12 - 0.57
Frac. sp3 C 0.02 0.13 0.08 0.57 -
Table A2: Maximum values of molecular descriptors in the pretraining set and designs of supervised finetuning and molecular task arithmetic for each setup. Highest values are displayed in boldface.
Task Setup Index Pretraining Finetuning Molecular Task Arithmetic
Frac. sp3C 0 1.00 1.00 1.00
1 1.00 1.00
2 1.00 1.00
3 1.00 1.00
4 1.00 1.00
logP 0 14.79 11.45 17.58
1 9.25 14.85
2 11.86 14.80
3 12.87 16.32
4 9.69 15.34
No. H Donors 0 16 9 15
1 9 15
2 11 14
3 10 15
4 9 15
No. Rings 0 14 10 16
1 11 17
2 12 20
3 11 14
4 10 17
TPSA 0 356.28 217.24 431.13
1 252.19 431.79
2 230.99 408.87
3 240.80 413.41
4 258.06 433.85
Positive Negative
Protein Training Validation Training Validation
F2 140 47 3208 1070
NR3C1 90 31 525 176
PPARD 12 5 570 190
Table A3: Dataset sizes used for bioactive molecule design experiments.
@@
Table A6: Number of successful design clusters, success rate, and number of clusters for reinforcement learning experiments with five different seeds. Average and standard deviation of each metric across runs are reported with 100,000 designs each.
Task Method No. Successful Design Clusters Success Rate No. Clusters
Low TPSA Pretraining 7542 $\pm$ 696 12% $\pm$ 1% 36245 $\pm$ 1867
MTA 22459 $\pm$ 912 39% $\pm$ 2% 32419 $\pm$ 1499
High TPSA Pretraining 1339 $\pm$ 137 1% $\pm$ 0% 43620 $\pm$ 1197
MTA 36492 $\pm$ 851 48% $\pm$ 1% 42304 $\pm$ 586
@@
Fig. A3: Docking score distribution of 100,000 pretraining molecules for the studied proteins.
@@
Fig. A6: Side effects of transfer learning strategies. Kolmogorov-Smirnov distance between the molecular properties of the designs on SMILES tasks and the pretraining set is computed for off-target properties. Lower distance to the pretraining set indicates that the conditioning caused less change in non-conditioned molecular properties. Bar heights display the means across five training splits, and error bars display the standard deviations.
Fig. A7: Side effects of conditioning strategies and pretraining. Compuations are described in (Fig. A6). Bar heights display the means across five training splits, and error bars display the standard deviations. In 32 of 40 comparisons, molecular task arithmetic caused less side-effects.
@@
Fig. A10: Zero-shot dual objective molecule design on canonical SMILES. Models are trained on randomized SMILES representation with sequential supervised finetuning and molecular task arithmetic to design molecules that possess two task properties simultaneously. Number of clusters and success rates are measured. Experiments are repeated five times, and the mean (bar height) and standard deviation (error bar) are reported.
[…]
@@
Fig. A14: Structural similarity distributions of top-1000 scoring molecules for finetuning and molecular task arithmetic in few-shot settings. Structural similarity is computed as Tanimoto similarity of extended connectivity fingerprints (radius=2, nBits=2048).
[…]